OpenStudy (anonymous):

find equation for line perpendicular to the line y=-4x-3; containing the point 3,3 ; final answer in slope intercept form

OpenStudy (anonymous):

anybody?

OpenStudy (anonymous):

y=(1/4)x+9/4

OpenStudy (anonymous):

i know the answer, its how u get there i dont

OpenStudy (anonymous):

if slope of two perpendicular lines follow the relation. m1*m2=-1 Use it to get slope as 1/4 write equation as y=1/4x+c find c by putting 3,3 in equation

OpenStudy (anonymous):

you dont neede to use point slope form?

OpenStudy (anonymous):

if the line of the given points are perpinduclar, then the slope will be the opposite recibrical. what you have to do next is simply use 1/4, and points 3,3 and plug it into y=mx+b what you have is 3=1/4*3+b. then you solve for b, which is simple algebra. after that, you just plug it all into y=mx+b and get y=1/4x+9/4

OpenStudy (anonymous):

can you use point slope form?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

how would that work?

OpenStudy (anonymous):

y-3=1/4(x-3)

OpenStudy (anonymous):

i appreciate it!!

OpenStudy (anonymous):

I really hope im right

OpenStudy (anonymous):

you r- i got the answer key

OpenStudy (anonymous):

sweeeet